# Spectral Bias Mitigation via xLSTM-PINN: Memory-Gated Representation Refinement for Physics-Informed Learning

### v1

@@ Abstract
Physics-informed learning for PDEs is surging across scientific computing and industrial simulation, yet prevailing methods face spectral bias, residual–data imbalance, and weak extrapolation. We introduce a representation-level spectral-remodeling xLSTM-PINN that combines gated-memory multiscale feature extraction with adaptive residual–data weighting to curb spectral bias and strengthen extrapolation. Across four benchmarks, we integrate gated cross-scale memory, a staged frequency curriculum, and adaptive residual reweighting, and verify with analytic references and extrapolation tests, achieving markedly lower spectral error and RMSE and a broader stable learning-rate window. Frequency-domain benchmarks show raised high-frequency kernel weights and a right-shifted resolvable bandwidth, shorter high-k error decay and time-to-threshold, and narrower error bands with lower MSE, RMSE, MAE, and MaxAE. Compared with the baseline PINN, we reduce MSE, RMSE, MAE, and MaxAE across all four benchmarks and deliver cleaner boundary transitions with attenuated high-frequency ripples in both frequency and field maps. This work suppresses spectral bias, widens the resolvable band and shortens the high-k time-to-threshold under the same budget, and—without altering AD or physics losses—improves accuracy, reproducibility, and transferability.
@@
{highlights}
xLSTM-PINN reshapes spectra via memory gating and residual micro-steps.
Suppresses spectral bias: lifts high-frequency gain and widens resolvable band.
Enhances high-frequency learnability without altering loss or differentiation paths. {graphicalabstract} [Uncaptioned image]
@@ 1 Introduction
Physics-Informed Neural Networks (PINNs)[raissi2019physics, xing2025modeling, li20252d, xu2025preprocessing, lin2025monte, yokota2024physics] have emerged as a powerful framework for solving partial differential equations (PDEs) by embedding physical laws directly into neural training. However, their effectiveness is fundamentally limited by a pronounced spectral bias—the inherent tendency of neural networks to learn low-frequency components faster than high-frequency modes. This imbalance impedes the accurate resolution of steep gradients, sharp interfaces, and high-wavenumber dynamics, which are crucial in many real-world physical systems. Overcoming this spectral bias is therefore central to enhancing the fidelity and applicability of physics-informed learning.
Numerous studies have attempted to mitigate spectral bias in PINNs by modifying network architectures, introducing Fourier-based features[hou2025fourier, aghaee2024performance, ortiz2025physics, li2024solving, sallam2023use], or designing frequency-aware sampling strategies[zeng2025application, yang2023dmis, daw2022mitigating, celaya2025adaptive]. While these efforts have yielded partial success—such as enhanced recovery of oscillatory solutions or localized accuracy improvement—they often incur trade-offs in scalability, stability, or implementation complexity. Some approaches[farhani2022momentum, fesser2023understanding, mustajab2024physics, liu2025diminishing] require manual frequency supervision or spectral priors, while others tightly couple the training process to specific PDE types or spatial domains, thus undermining generalization. Moreover, most existing schemes do not directly address the representational inefficiency at the layer level that underlies spectral degradation, especially in deeper PINN stacks[kashefi2025pointnet, xu2024building, kashefi2025physics, kashefi2024kolmogorov, kashefi2022physics, tao2025operator]. This motivates the exploration of architectural innovations that can implicitly enhance spectrum resolution without altering loss terms, gradient flows, or optimization schedules.
Among recent advances in neural representation design, xLSTM[beck2024xlstm] has emerged as a promising architecture featuring gated memory dynamics and residual micro-steps within each layer. This structure introduces intra-layer recursion that enables richer, more flexible information flow without increasing the number of trainable parameters. By repeatedly refining intermediate states through lightweight memory updates, xLSTM effectively deepens the computational trajectory of each layer and shows strong potential in capturing high-frequency patterns that conventional feedforward models often miss. Although originally developed for general sequence modeling, its capacity for dynamic, spectrum-sensitive representation makes it a compelling candidate for addressing the spectral bias inherent in PINNs.
In this work, we embed xLSTM blocks into the representation layers of PINNs, introducing memory gating and residual micro-steps without altering the physics loss functions or automatic differentiation pathways. This minimal yet principled modification reshapes the spectrum of the underlying neural tangent kernel, systematically lifting high-frequency eigenmodes and expanding the resolvable bandwidth. Through both theoretical analysis and frequency-domain benchmarks, we demonstrate that xLSTM-PINN significantly mitigates spectral bias and accelerates convergence along high-wavenumber directions. Numerical experiments across diverse PDE settings further confirm consistent improvements in accuracy and error reduction. These results establish xLSTM-PINN as a robust, architecture-level remedy to longstanding spectral limitations in physics-informed learning.
@@ 2.1 Architecture and Operator Mapping: Memory-Gated Residual Micro-Steps in xLSTM-PINN
We define the physical domain $\Omega\subset\mathbb{R}^{d}$ and feed the coordinate $\bf{x}$ into a representation network composed of “gated memory–residual micro-steps,” and then use a linear head to output a scalar/vector field $u_{\theta}(\bf{x})$. The network departs from a conventional PINN only at the representation level; we keep automatic differentiation (AD) and the construction of physics losses identical. Fig. 1 illustrates the closed loop “representation network $\to$ AD $\to$ physics loss $\to$ optimization.”
We write the mapping as
$u_{\theta}(\bf{x})=W_{\text{out}}\,u^{(L)}(\bf{x})+b_{\text{out}},\hskip 22.99988ptu^{(0)}(\bf{x})=\phi\!\left(W_{\text{in}}\bf{x}+b_{\text{in}}\right),$ (1)
where $\phi$ denotes $\tanh$, and $L$ denotes the number of blocks. Each layer applies one xLSTM block $\mathcal{B}^{(\ell)}$ to transform $u^{(\ell)}$ into $u^{(\ell+1)}$
@@
An xLSTM block introduces an explicit “micro-time” iteration and gated memory inside a single layer. For an input $u\in\mathbb{R}^{W}$ with channel width $W$, we set $(h_{0},c_{0},n_{0},m_{0})=(0,0,0,0)$ and, for $t=0,\dots,S-1$, compute
@@
We then apply a stabilization re-scaling to prevent numerical blow-up/vanishing of accumulated memory (proof and invariance appear in Sec I of Supplementary Materials (SM))
@@
We update the binary state of “memory–duty cycle” and produce the gated output
@@
and refine the representation via a residual micro-step
@@
where $\psi$ denotes $\tanh$, and $P\in\mathbb{R}^{W\times W}$. Eqs. (1)-(6) instantiate, at the channel level, the ring-shaped memory path (constant-error carousel), the three gates/exponential gate, and the residual merge $\oplus$ in the “xLSTM Blocks” of Fig. 1. We amplify the “effective computational depth” of a single block by executing $S$ micro-steps with shared parameters, and thus enhance representational granularity without increasing width. The block outputs $u_{S}$.
Figure 1: xLSTM-PINN overview and intra-block recursion. We embed coordinates by a linear–tanh map to $u^{(0)}$, apply per-layer xLSTM blocks with $S$ micro-steps, and output $u_{\theta}$ via a linear head; AD provides PDE residuals and boundary terms for the loss.
@@
To strengthen nonlinear coupling across channels while preserving stable training, we attach a gated feedforward mixer to the tail of the block
@@
where $\varphi_{1,2}$ denote $\tanh$, and $\gamma$ uses $\sigma$ as a channel gate. We then optionally apply layer normalization (Sec II of Supplementary Materials (SM) analyzes its effect on gradient flow). We define
@@
as the inter-layer composite mapping, so that each layer repeats “gated memory Eqs.(1)–(6) $\to$ gated feedforward mixing Eq.(7) $\to$ linear–nonlinear shaping Eq.(8).” The parameter count of the whole network scales, to leading order, as
@@
We embed the above representation network into a PINN by constructing physics terms via AD on sampling sets $\mathcal{S}_{r},\mathcal{S}_{D},\mathcal{S}_{N},\mathcal{S}_{\text{IC}}$. Let $\mathcal{L}$ denote a generic operator (including coefficients, sources, and coordinate transforms), and let $\bf{n}$ denote the outward normal. The overall objective reads
@@
which aligns with the “representation network $\to$ AD $\to$ multi-term physics losses $\to$ weighted sum” shown in Fig. 1 (upper panel). We minimize $\mathcal{J}$ using a first-order optimizer (e.g., Adam[kingma2014adam]), and we confine all “spectral-bias–related” modifications strictly to the gating and micro-step mechanisms of the representation $\{\mathcal{B}^{(\ell)}\}$, so that we can compare learning dynamics across representations under identical sampling, optimization protocol, and parameter budget.
At the theoretical level, we write linearized training dynamics under the Neural Tangent Kernel (NTK) approximation as modal decay $c_{j}(t)\approx\mathrm{e}^{-\eta\lambda_{j}t}c_{j}(0)$. In Sec 2.2 we prove that when the residual micro-steps and gated memory inside a block flatten the tail of the empirical kernel’s eigenvalue distribution over the high-frequency subspace, the exponential decay rates of high-wavenumber modes increase uniformly; under a fixed budget this shift pushes the resolvable bandwidth $k^{\star}(\varepsilon)$ to the right (derivations and sufficient conditions appear in Sec III of Supplementary Materials Supplementary Materials (SM)). This conclusion matches Fig. 1 one-to-one: the blue “xLSTM Blocks” perform reparameterization and provide memory channels, while the purple frame preserves the AD–physics constraint pathway of PINNs.
We thus cleanly decouple “network architecture” from “physics constraints”: Fig. 1 replaces only the representation side while keeping the loss and optimization sides intact; Eqs. (1)–(8) give reproducible hierarchical equations, and Eq. (10) embeds them into a standard PINN objective. In Sec 2.2 we will characterize the resulting differences in convergence structure in spectral coordinates and validate them with independent frequency-domain benchmarks.
@@ 2.2 Spectral-Bias Suppression and Frequency-Domain Gain: Kernel-Tail Lifting in xLSTM-PINN
We define spectral bias as the imbalance of convergence weights across frequency modes. Let $\Omega=[-1,1]^{d}$ with probability measure $\mu$. Using the orthogonal basis
@@
we denote the model output $u_{\theta}(\mathbf{x},t)$ and define the frequency error coefficient
@@
Under the Neural Tangent Kernel (NTK) linearization, we write the training dynamics as
@@
where the kernel operator $K$ admits, with respect to $\mu$, the eigendecomposition $K\phi_{\mathbf{k}}=\lambda(\mathbf{k})\,\phi_{\mathbf{k}}$. Projecting Eq. (13) onto $\phi_{\mathbf{k}}$ yields the scalar ODE
@@
The decay of $\lambda(\mathbf{k})$ with $|\mathbf{k}|$ provides the mathematical characterization of spectral bias.
Figure 2: Frequency-domain benchmark evidences spectral-bias suppression and bandwidth expansion of xLSTM-PINN. Under matched training budgets and model size, we probe the spectrum with plane waves and report endpoint error $E_{T}(|k|)$, gain $G(|k|)=E_{\text{base}}/E_{\text{xLSTM}}$, and time-to-threshold $\tau(|k|)$, with uncertainty shown as confidence bands and the resolution frontier $k^{*}(\varepsilon)$ indicating the learnable bandwidth. This benchmark isolates representation effects and directly quantifies improved high-frequency learnability driven by memory gating and residual micro-steps.
(a) Endpoint error vs frequency: $E_{T}(|k|)$ decreases systematically in the mid-high range, with the error plateau lowered and extended rightward.
(b) Spectral gain vs frequency: $G(|k|)$ exceeds 1 over a broad band; non-overlapping confidence intervals highlight robust, high-frequency advantages.
(c) Time-to-threshold vs frequency: $\tau(|k|)$ shifts downward with a right-shifted $k^{*}(\varepsilon)$, demonstrating higher resolvable wavenumbers at the same budget, reduced spectral bias, and improved accuracy.
We introduce the xLSTM block in Fig. 1 on the representation side and perform $S$ gated micro-steps within one layer:
@@
where $\psi=\tanh$. We write the explicit form of $\mathrm{LSTM}(\cdot)$ as
@@
We linearize the $S$ micro-steps of one block with respect to the input representation $u$. Assume $\Delta u_{t}\approx A\,u_{t}$ with
@@
where $J_{u}h$ denotes the Jacobian of $h$ with respect to $u$. Using Eq. (16), we obtain
@@
with
@@
At the stable initialization $(h_{t},c_{t})=(0,0)$, Eq.  (18) gives an explicit $J_{u}h$. Hence a single block admits the effective linear mapping
@@
Let $\Phi\in\mathbb{R}^{N\times W}$ denote the baseline feature matrix of one layer. The corresponding kernels read
@@
Let $\Phi^{\top}\phi_{\mathbf{k}}=v_{\mathbf{k}}\in\mathbb{R}^{W}$. Then
@@
Let $B:=(A+A^{\top})/2$ denote the symmetric part. For any $v$ we have
@@
where $\rho_{B}(v):=\dfrac{v^{\top}Bv}{\|v\|_{2}^{2}}\geq 0$. From Eqs. (22)–(23) we derive the bound
@@
We introduce the frequency gain
@@
If $\alpha(\mathbf{k})$ increases monotonically with $|\mathbf{k}|$, the kernel spectrum raises its high-frequency tail monotonically. Substituting Eqs. (24)–(25) into the closed-form solution of the frequency error yields three direct conclusions:

### v2

# xLSTM-PINN: Memory-Gated Spectral Remodeling for Physics-Informed Learning

@@
[orcid=0009-0004-0202-3641] [orcid=0009-0003-6352-6317] [orcid=0000-0002-8573-450X] [orcid=0009-0003-7880-0235] [orcid=0009-0001-6432-2733] 1]organization=Nanophotonics and Biophotonics Key Laboratory of Jilin Province, School of Physics, Changchun University of Science and Technology, city=Changchun, postcode=130022, country=P.R. China
xLSTM-PINN: Memory-Gated Spectral Remodeling for Physics-Informed Learning
Ze Tao    Darui Zhao    Fujun Liu fjliu@cust.edu.cn    Ke Xu    Xiangsheng Hu [
@@ A B S T R A C T
Physics-informed neural networks (PINN) face significant challenges from spectral bias, which impedes their ability to model high-frequency phenomena and limits extrapolation performance. To address this, we introduce xLSTM-PINN, a novel architecture that performs representation-level spectral remodeling through memory gating and residual micro-steps. Our method consistently achieves markedly lower spectral error and root mean square error (RMSE) across four diverse partial differential equation (PDE) benchmarks, along withhhh a broader stable learning-rate window. Frequency-domain analysis confirms that xLSTM-PINN elevates high-frequency kernel weights, shifts the resolvable bandwidth rightward, and shortens the convergence time for high-wavenumber components. Without modifying automatic differentiation or physics loss constraints, this work provides a robust pathway to suppress spectral bias, thereby improving accuracy, reproducibility, and transferability in physics-informed learning.
@@ 1 Introduction
Physics-Informed Neural Networks (PINNs) [raissi2019physics, xing2025modeling, li20252d, xu2025preprocessing, lin2025monte, yokota2024physics] have established themselves as a powerful framework for solving partial differential equations (PDEs) by directly embedding physical laws into the learning process. Despite their conceptual appeal, these models face a fundamental limitation known as spectral bias—a systematic tendency to prioritize low-frequency components during training while struggling to capture high-frequency modes. This imbalance manifests as poor resolution of steep gradients, blurred interfaces, and inaccurate high-wavenumber dynamics, significantly restricting PINNs’ practical utility across many scientific and engineering domains. Consequently, overcoming spectral bias has become a central challenge in advancing physics-informed learning.
In response to this challenge, researchers have developed various mitigation strategies that primarily focus on architectural modifications, Fourier-based feature embeddings [hou2025fourier, aghaee2024performance, ortiz2025physics, li2024solving, sallam2023use], and frequency-aware sampling schemes [zeng2025application, yang2023dmis, daw2022mitigating, celaya2025adaptive]. Although these approaches have demonstrated improved performance on oscillatory problems, they frequently introduce practical limitations regarding scalability, numerical stability, or implementation complexity. Techniques requiring manual frequency supervision [farhani2022momentum, fesser2023understanding, mustajab2024physics, liu2025diminishing] or domain-specific spectral priors further constrain generalization potential, while most methods fail to address the core issue of representational inefficiency within individual network layers [kashefi2025pointnet, xu2024building, kashefi2025physics, kashefi2024kolmogorov, kashefi2022physics, tao2025operator]. This landscape motivates the development of architectures that intrinsically enhance spectral resolution without modifying loss formulations or optimization procedures.
The recently introduced xLSTM architecture [beck2024xlstm] presents a promising direction through its novel combination of gated memory mechanisms and residual micro-steps implemented at each layer. By incorporating intra-layer recursion, xLSTM enables more dynamic information flow without increasing parameter counts, effectively deepening computational paths through lightweight memory operations. This design exhibits particular strength in capturing high-frequency patterns that typically challenge conventional feedforward networks, suggesting its potential applicability to spectral bias challenges in physics-informed learning. Originally designed for sequence modeling, xLSTM’s capacity for spectrum-sensitive representation offers an intriguing foundation for PINN enhancements.
Building upon these insights, we introduce xLSTM-PINN, which integrates xLSTM blocks as core representation components while maintaining standard physics-constrained learning protocols. Our approach preserves automatic differentiation pathways and physics loss formulations unchanged, focusing exclusively on architectural improvements to reshape the neural tangent kernel’s spectral properties. Through theoretical analysis and empirical validation, we demonstrate that this integration systematically elevates high-frequency eigenmodes and extends resolvable bandwidth. Extensive experiments across diverse PDE benchmarks confirm that xLSTM-PINN achieves accelerated convergence along high-wavenumber directions and consistent error reduction, establishing it as an effective architecture-level solution to the persistent challenge of spectral bias in physics-informed learning.
@@ 2.1 Architecture and Operator Mapping: Memory-Gated Residual Micro-Steps in xLSTM-PINN
We begin by defining the physical domain $\Omega\subset\mathbb{R}^{d}$ and constructing a representation network that processes coordinate inputs $\mathbf{x}$ through gated memory-residual micro-steps, ultimately producing a scalar or vector field $u_{\theta}(\mathbf{x})$ via a linear output layer. This architecture modifies conventional PINNs exclusively at the representation level while preserving identical automatic differentiation procedures and physics loss constructions. The complete workflow follows the following sequence: representation network $\rightarrow$ automatic differentiation $\rightarrow$ physics loss computation $\rightarrow$ optimization, as illustrated in Fig. 1.
The network mapping is formally defined as:
$\bm{u}_{\theta}(\bf{x})=W_{\text{out}}\,u^{(L)}(\bf{x})+b_{\text{out}},\hskip 22.99988ptu^{(0)}(\bf{x})=\phi\!\left(W_{\text{in}}\bf{x}+b_{\text{in}}\right),$ (1)
where $\phi$ denotes the $\tanh$ activation function, and $L$ represents the total number of processing blocks. Each layer transforms its input through an xLSTM block $\mathcal{B}^{(\ell)}$ according to:
@@
The xLSTM block introduces sophisticated memory mechanisms through micro-time iterations within each layer. For an input $\mathbf{u}\in\mathbb{R}^{W}$ with channel width $W$, we initialize states $(h_{0},c_{0},n_{0},m_{0})=(0,0,0,0)$ and compute for $t=0,\dots,S-1$:
@@
To ensure numerical stability and prevent memory explosion or vanishing gradients (proof and invariance appear in Sec I of Supplementary Materials (SM)), we apply stabilization scaling:
@@
The memory states are then updated through gating operations:
@@
followed by residual refinement of the representation:
@@
where $\psi$ denotes $\tanh$, and $P\in\mathbb{R}^{W\times W}$. Collectively, Eqs. (1)-(6) implement the ring-shaped memory path, gating mechanisms, and residual connections depicted in the xLSTM blocks of Fig. 1. This design amplifies effective computational depth through $S$ micro-steps with shared parameters, enhancing representational capacity without increasing network width.
Figure 1: xLSTM-PINN overview and intra-block recursion. We embed coordinates by a linear–tanh map to $u^{(0)}$, apply per-layer xLSTM blocks with $S$ micro-steps, and output $u_{\theta}$ via a linear head; Automatic Differentiation (AD) provides PDE residuals and boundary terms for the loss.
@@
To further strengthen cross-channel interactions while maintaining training stability, we incorporate a gated feedforward mixer:
@@
where $\varphi_{1,2}$ represent $\tanh$ activations, and $\gamma$ employs $\sigma$ as a channel-wise gating function. Optional layer normalization (analyzed in Sec II of Supplementary Materials (SM)) may be applied to improve gradient flow. The complete block transformation is defined as:
@@
establishing the repeating sequence: gated memory (Eqs. (1)–(6)) $\rightarrow$ feedforward mixing (Eq. (7)) $\rightarrow$ linear-nonlinear projection (Eq. (8)). The resulting parameter complexity scales as:
@@
This representation network integrates into the PINN framework through physics-constrained optimization on sampling sets $\mathcal{S}_{r},\mathcal{S}_{D},\mathcal{S}_{N},\mathcal{S}_{\text{IC}}$. For a differential operator $\mathcal{L}$ (incorporating coefficients, sources, and coordinate transformations) and an outward normal $\mathbf{n}$, the composite objective function becomes:
@@
which maintains the standard PINN optimization pathway while concentrating spectral-bias modifications within the representation components $\{\mathcal{B}^{(\ell)}\}$. We minimize $\mathcal{J}$ using first-order methods (e.g., Adam [kingma2014adam]), enabling fair comparison of learning dynamics across architectures under consistent sampling, optimization, and parameter budgets.
Theoretical analysis under the Neural Tangent Kernel framework reveals that linearized training dynamics follow modal decay $c_{j}(t)\approx\mathrm{e}^{-\eta\lambda_{j}t}c_{j}(0)$. As established in Sec 2.2, the residual micro-steps and gated memory collectively flatten the eigenvalue distribution tail in high-frequency subspaces, uniformly accelerating exponential decay rates for high-wavenumber modes. This spectral reshaping extends the resolvable bandwidth $k^{\star}(\varepsilon)$ under fixed computational resources, with detailed derivations provided in Sec III of the Supplementary Materials (SM). This theoretical insight aligns precisely with the architectural shown in Fig. 1, where xLSTM blocks handle representation refinement while the standard PINN framework preserves automatic differentiation and physical constraints.
This design achieves clean separation between network architecture and physical constraints: the representation pathway undergoes comprehensive modification through Eqs. (1)–(8), while the loss formulation (Eq. (10)) and optimization procedures remain consistent with conventional PINNs. The subsequent section characterizes the resulting convergence improvements in spectral coordinates and validates them through independent frequency-domain benchmarks.
@@ 2.2 Spectral-Bias Suppression and Frequency-Domain Gain: Kernel-Tail Lifting in xLSTM-PINN
We begin by formally characterizing spectral bias as the systematic imbalance in convergence rates across different frequency modes during neural network training. To establish a rigorous mathematical framework, consider the domain $\Omega=[-1,1]^{d}$ equipped with probability measure $\mu$, and define the orthogonal basis functions:
@@
which enables spectral decomposition of both the target function and model outputs. The frequency-dependent error coefficient captures the projection of the approximation error onto each spectral mode:
@@
Within the Neural Tangent Kernel theoretical framework, the training dynamics follow the integral evolution:
@@
where the kernel operator $K$ admits the eigendecomposition $K\phi_{\mathbf{k}}=\lambda(\mathbf{k})\,\phi_{\mathbf{k}}$ with respect to $\mu$. Projecting this dynamics onto individual spectral modes yields decoupled ordinary differential equations:
@@
The central manifestation of spectral bias emerges from the characteristic decay of $\lambda(\mathbf{k})$ with increasing $|\mathbf{k}|$, which systematically slows high-frequency convergence.
Figure 2: Frequency-domain benchmark evidences spectral-bias suppression and bandwidth expansion of xLSTM-PINN. Under matched training budgets and model size, we probe the spectrum with plane waves and report endpoint error $E_{T}(|k|)$, gain $G(|k|)=E_{\text{base}}/E_{\text{xLSTM}}$, and time-to-threshold $\tau(|k|)$, with uncertainty shown as confidence bands and the resolution frontier $k^{*}(\varepsilon)$ indicating the learnable bandwidth. This benchmark isolates representation effects and directly quantifies improved high-frequency learnability driven by memory gating and residual micro-steps. (a) Endpoint error vs frequency: $E_{T}(|k|)$ decreases systematically in the mid-high range, with the error plateau lowered and extended rightward; (b) Spectral gain vs frequency: $G(|k|)$ exceeds 1 over a broad band; non-overlapping confidence intervals highlight robust, high-frequency advantages; (c) Time-to-threshold vs frequency: $\tau(|k|)$ shifts downward with a right-shifted $k^{*}(\varepsilon)$, demonstrating higher resolvable wavenumbers at the same budget, reduced spectral bias, and improved accuracy.
The xLSTM architecture introduces transformative modifications to address this spectral limitation through integrated gated memory operations and residual connections. Each layer executes $S$ sequential micro-steps following the recurrence relation:
@@
where $\psi=\tanh$ providing nonlinear refinement. The LSTM cell implements precise memory control through gating mechanisms:
@@
To analytically quantify the spectral impact of these architectural innovations, we linearize the micro-step sequence around the initialization point. Assuming local linear behavior $\Delta u_{t}\approx A\,u_{t}$, the transformation Jacobian becomes:
@@
where $J_{u}h$ denotes the Jacobian of $h$ with respect to $u$. Through systematic differentiation of Eq. (16), we obtain the explicit Jacobian expression:
@@
with diagonal matrices capturing gate activation derivatives:
@@
At the stable initialization $(h_{t},c_{t})=(0,0)$, Eq. (18) provides a computable $J_{u}h$, leading to the effective linear mapping after $S$ micro-steps:
@@
This linearized analysis reveals the spectral enhancement mechanism. For baseline features $\Phi\in\mathbb{R}^{N\times W}$, the corresponding kernels read:
@@
Projecting onto spectral modes via $\Phi^{\top}\phi_{\mathbf{k}}=v_{\mathbf{k}}\in\mathbb{R}^{W}$ yields the critical eigenvalue relations:
@@
The spectral enhancement mechanism becomes mathematically explicit through eigenvalue ratio analysis. Defining the symmetric component $B:=(A+A^{\top})/2$, we establish the foundational inequality:
@@
where $\rho_{B}(v):=\dfrac{v^{\top}Bv}{\|v\|_{2}^{2}}\geq 0$ represents the Rayleigh quotient. From Eqs. (22)–(23) we derive the fundamental lower bound:
@@
Introducing the frequency gain function:
@@
we identify the precise mechanism for spectral bias suppression: monotonic increase of $\alpha(\mathbf{k})$ with $|\mathbf{k}|$ directly elevates the high-frequency tail of the kernel spectrum.
Substituting these analytical results into the frequency error dynamics produces three mathematically rigorous conclusions. The endpoint error ratio:
@@
quantifies exponential improvement in final accuracy. The convergence time ratio:
@@
demonstrates accelerated high-frequency learning. The resolvable bandwidth condition: